OpenStudy (anonymous):

Consider the polynomial function f defined by f(x) = x^3 + 2x − 8 Let the inverse of f be g.Find the best linear approximation to g when the argument of g is close to -5.

OpenStudy (anonymous):

I've calculated that derivative of g at -5 would be 1/5. What do I do next?

OpenStudy (anonymous):

\[g \prime(x)=\frac{ 1 }{ f \prime (g(x)) }\]

OpenStudy (solomonzelman):

An example problem to your question. I hope that will address your question, emcrazy14\(\small : ) \) \(\color{#0000ff }{\huge \displaystyle_{_{_\text{_________________________________}}} }\) \(\color{#0000ff }{ \displaystyle \bf Problem: }\) Consider the (polynomial) function \(\color{#000000 }{ \displaystyle f}\) defined by\({\tiny\\[0.5em]}\) \(\color{#000000 }{ \displaystyle f(x)=x^3-4x^2+10 }\), and let \(\color{#000000 }{ \displaystyle f^{-1}(x)=g(x) }\)\({\tiny\\[0.5em]}\) Find the best linear approximation to \(\color{#000000 }{ \displaystyle g}\), near\({\tiny\\[0.5em]}\) the point \(\color{#000000 }{ \displaystyle (3.45,~3.45) }\). \(\color{#0000ff }{\huge \displaystyle_{_{_\text{_________________________________}}} }\) \(\color{#0000ff }{ \displaystyle \bf Solution: }\) \(\color{red}{_■}\) Step 1\( : \) Find the inverse of \(\color{#000000 }{ \displaystyle f }\). \(\color{#000000 }{ \displaystyle f(x)=x^3-4x^2+10 }\) I am putting \(\color{#000000 }{ \displaystyle y }\) instead of \(\color{#000000 }{ \displaystyle f(x) }\) for convenience. \(\color{#000000 }{ \displaystyle y=x^3-4x^2+10 }\) Then, I swap the x and y (with each other). \(\color{#000000 }{ \displaystyle x=y^3-4y^2+10 }\) Solve this for \(\color{#000000 }{ \displaystyle y }\) is painful, be have already found the inverse technically. So we just have to remember that really \(\color{#000000 }{ \displaystyle f^{-1}(x)=g(x)=y }\) and we need to approximate this linearly as at \(\color{#000000 }{ \displaystyle x=3.45 }\) So we will leave it as, \(\color{#000000 }{ \displaystyle x=y^3-4y^2+10 }\) and use implicit differentiation instead. \(\color{red}{_■}\) Step 2\( : \) Differentiate \(\color{#000000 }{ \displaystyle x=y^3-4y^2+10 }\). CHAIN RULE NOTE\( : \) Note, that the chain rule of "y-prime" is applied. Just like the derivative of \((x^2+3)^{100}\) is NOT just \(100(x^2+3)^{99}\), RATHER, you differentiate the part inside as well and multiply (via Chain Rule) and so you would get \(100(x^3+3)^{99}\color{blue}{\times 3x^2 }\). The blue comes from the "Chain Rule". That simplifies to \(300x^2(x^3+3)^{99}\), but that is NOT the point. THE POINT IS: when you differentiate \(y\), you also multiply times the derivative of y (times y'), since y is also a function of x. Therefore, when you differentiate (for instance) \(y^4\) (with respect to x), you get, \(4y^3\color{blue}{\times y'}\). DIFFERENTIATING the function \(\color{#000000 }{ \displaystyle x=y^3-4y^2+10 }\) \( : \) \(\color{#000000 }{ \displaystyle 1=3y^2y'-8yy'+0 }\) Then, simplifying and solving for y', \(\color{#000000 }{ \displaystyle 1=3y^2y'-8yy' }\) \(\color{#000000 }{ \displaystyle 1=(3y^2-8y)y' }\) \(\color{#000000 }{ \displaystyle y'=1/(3y^2-8y) }\) Just in case y' notation confuses you, that is also an equivalent to saying \(\color{#000000 }{ \displaystyle \frac{dy}{dx}=\frac{1}{3y^2-8y} }\) \(\color{red}{_■}\) Step 3\( : \) Find slope at \((3.45, 3.45)\). To find the slope of the function at the point (3.45, 3.45), We plug in x=3.45 and y=3.45 (There isn't an "x" in the derivative, so we just plug y=3.45) \(\color{#000000 }{ \displaystyle g'(3.45,3.45)=\frac{1}{3(3.45)^2-8(3.45)} =\frac{1}{8.1075}\approx 0.123 }\) And that is our slope. \(\color{red}{_■}\) Step 4\( : \) Linearization of \(g\) at \(\color{#000000 }{ \displaystyle (3.45,~3.45) }\). Given the point (3.45, 3.45)\((3.45, 3.45)\) and slope \(m=0.123\); y-3.45=0.123(x-3.45)\(y-3.45=0.123(x-3.45)\) CHECKING: (In the graph I entered the EXACT slope) https://www.desmos.com/calculator/3sdq7ik3df